# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ Appendix C Detailed proof
From Nesterov and Polyak (2006, Theorem 1), we have
$\min_{k=1,...,K_{1}}\nu_{\overline{\overline{L}}_{g}}(y_{k})\leq\frac{8}{3}\left(\frac{3(g(x,y_{0})-\underline{g})}{2K_{1}\overline{\overline{L}}_{g}}\right)^{1/3},$
where
$\nu_{\overline{\overline{L}}_{g}}(y_{k})=\max\left\{\sqrt{\frac{1}{\overline{\overline{L}}_{g}}\|\nabla_{y}g(x,y_{k})\|},-\frac{2}{3\overline{\overline{L}}_{g}}\lambda_{n}(\nabla^{2}_{yy}g(x,y_{k}))\right\}.$
Without loss of generality, we assume $\arg\min_{k=1...,K_{1}}\nu_{\overline{\overline{L}}_{g}}(y_{k})$ exactly equals $K_{1}$ which is defined in (4.4). It follows
$\|\nabla_{y}g(x,y_{K_{1}})\|$ $\leq\overline{\overline{L}}_{g}\left(\nu_{\overline{\overline{L}}_{g}}(y_{K_{1}})\right)^{2}$
$\leq\overline{\overline{L}}_{g}\left(\frac{8}{3}\left(\frac{3(g(x,y_{0})-\underline{g})}{2K_{1}\overline{\overline{L}}_{g}}\right)^{1/3}\right)^{2}$
$<\frac{64\overline{\overline{L}}_{g}}{9}\left(\frac{3(\overline{g}_{0}-\underline{g})}{2\frac{256(\overline{g}_{0}-\underline{g})\left(\overline{\overline{L}}_{g}\right)^{1/2}}{9\left(\min\left\{\alpha(\delta,\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}}),\frac{(\mu(\delta))^{2}}{4\overline{\overline{L}}_{g}}\right\}\right)^{3/2}}\overline{\overline{L}}_{g}}\right)^{2/3}$
$=\min\left\{\alpha\left(\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}}\right),\frac{(\mu(\delta))^{2}}{4\overline{\overline{L}}_{g}}\right\},$ (C.2)
where $\overline{g}_{0}$ is from Proposition 3.1(1), and
$\lambda_{n}(\nabla^{2}_{yy}g(x,y_{K_{1}}))$ $\geq-\frac{3\overline{\overline{L}}_{g}}{2}\nu_{\overline{\overline{L}}_{g}}(y_{k})$
$\geq-4\overline{\overline{L}}_{g}\left(\frac{3(g(x,y_{0})-\underline{g})}{2K_{1}\overline{\overline{L}}_{g}}\right)^{1/3}$
$\geq-4\overline{\overline{L}}_{g}\left(\frac{3(\overline{g}_{0}-\underline{g})}{2\frac{768(\overline{g}_{0}-\underline{g})\overline{\overline{L}}^{2}_{g}}{\left(\mu(\delta)\right)^{3}}\overline{\overline{L}}_{g}}\right)^{1/3}$
$=-\frac{\mu(\delta)}{2},$ (C.3)
where $\lambda_{n}(\nabla^{2}_{yy}g(x,y_{K_{1}}))$ denotes the smallest eigenvalue of $\nabla^{2}_{yy}g(x,y_{K_{1}})$ and functions $\mu(\cdot)$ and $\alpha(\cdot,\cdot)$ are from Lemma 4.2 and Lemma 4.3. We next show that within $K_{1}$ iterations, the sequence enters the strongly convex neighborhood of a local minimizer, denoted by $y^{\text{cd}}(x)$, and eventually converges to this local minimizer.
Since $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$, $g(x,y)$ is a Morse function with respect to $y$. This implies that the set of stationary points is discrete and finite. Suppose the stationary point of $g(x,y)$ with respect to $y$ over $\mathcal{Y}^{\prime}$ is $p_{1}$, …, $p_{J}$. Consider balls $B(p_{j},\mu(\delta)/2\overline{\overline{L}}_{g})$. By Lemma 4.2, Lemma 4.3 and Lipschitz continuity of $\nabla^{2}_{yy}g(x,y)$ from Assumption 3.2(4), we know that for any $y\in\mathcal{Y}$ outside these balls, we have $\|\nabla_{y}g(x,y)\|\geq\alpha(\delta,\mu(\delta)/2\overline{\overline{L}}_{g})$, and for any $y$ in these balls, every eigenvalue $\lambda$ of $\nabla^{2}_{yy}g(x,y)$ satisfies $|\lambda|\geq\mu(\delta)-(\mu(\delta)/2\overline{\overline{L}}_{g})\cdot\overline{\overline{L}}_{g}=\mu(\delta)/2$. Note that, by Assumption 3.2(6) and 3.3, all iterates remain in $\mathcal{Y}$. Combining this with (C.9), we obtain that $y_{K_{1}}$ is in a ball $B(p_{j},\mu(\delta)/2\overline{\overline{L}}_{g})$ in which $g(x,y)$ is strongly convex with coefficient $\mu(\delta)/2$. That is, we have excluded from the set of candidates for $y_{K_{1}}$ the ball centered at those stationary points that are not local minima. We can also exclude balls $B(p_{i},\mu(\delta)/2\overline{\overline{L}}_{g})$ centered at local minima $p_{i}\in\mathcal{Y}^{\prime}\setminus\mathcal{Y}$. Since $p_{i}\notin\mathcal{Y}$ implies $g(x,p_{i})>g(x,y_{0})$, and $g(x,\cdot)$ is locally minimized at $p_{i}$, all points within $B(p_{i},\mu(\delta)/2\overline{\overline{L}}_{g})$ will also have function values strictly greater than $g(x,y_{0})$. Therefore, the iterates can never enter such neighborhoods.
Since $g(x,\cdot)$ is $\mu(\delta)/2$-strongly convex in
$B\!\left(p_{j},\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}}\right)$
and $p_{j}$ is a local minimizer, we have $\nabla_{y}g(x,p_{j})=0$. By the definition of strong convexity, for any $y$ in this ball,
$g(x,y)\geq g(x,p_{j})+\frac{\mu(\delta)}{4}\,\|y-p_{j}\|^{2}.$
In particular, for $y$ on the boundary of the ball, i.e., $\|y-p_{j}\|=\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}},$ we have
$g(x,y)\geq g(x,p_{j})+\frac{\mu(\delta)}{4}\left(\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}}\right)^{2}=g(x,p_{j})+\frac{(\mu(\delta))^{3}}{4(2\overline{\overline{L}}_{g})^{2}}.$
Therefore, the local level set $\mathcal{F}=\{y\in B(p_{j},\mu(\delta)/2\overline{\overline{L}}_{g}):g(x,y)-g(x,p_{j})\leq(\mu(\delta))^{3}/(16\overline{\overline{L}}_{g}^{2})\}$ is contained in the interior of the ball $B(p_{j},\mu(\delta)/2\overline{\overline{L}}_{g})$. Since $\|\nabla_{y}g(x,y_{K_{1}})\|<(\mu(\delta))^{2}/(4\overline{\overline{L}}_{g})$ can imply that $y_{K_{1}}$ lies in $\mathcal{F}$, by Nesterov and Polyak (2006, Lemma 2), all subsequent points lie in this level set. It’s not hard to see $y_{k}$ converge to $y^{\text{cd}}(x):=p_{j}$ as $k$ goes to infinity. Starting from $y_{K_{1}}$, we perform $K_{2}$ iterations of the cubic-regularized Newton method. Then, by applying Nesterov and Polyak (2006, Theorem 1) once again, as in (C.9), we obtain a point whose gradient norm is at most $\mu(\delta)\rho/2$. That is, among the iterates from the $K_{1}$-th to the $(K_{1}+K_{2})$-th step, there exists a point whose gradient norm is at most $\mu(\delta)\rho/2$. Denote this point by $\hat{y}(x)$. Note that the last $K_{2}$ iterations are all performed within the local level set $\mathcal{F}$, in which $g(x,\cdot)$ is $\mu-(\delta)/2$-strongly convex, then we have
$\|\hat{y}(x)-y^{\text{cd}}(x)\|\leq\rho.$
@@
Note that $K$ satisfies (4.4), according to Lemma 4.4, we have $\|\hat{y}(z)-y^{\text{cd}}(z)\|\leq\rho$ for any $z\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$. By Lipschitz continuity of $f$ (from Proposition 3.1(2)), it follows $|f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))|\leq L_{f}\rho$. From (3.8), we know that $|f(z,y^{\text{cd}}(z))|\leq\overline{f}$ for any $z\in\mathcal{X}$. By an argument similar to that of (3.8), we have $|f(z,\hat{y}(z))|\leq\overline{f}$ for any $z\in\mathcal{X}$ as well. Thus, the following holds:
@@
Note that we assume that the number of lower-level iterations satisfies $K\geq K_{0}$ in Assumption 3.3. $y^{\text{cd}}(x)$ is always in the level set $\{y:g(x,y)\leq g(x,y_{0})\}$. By Proposition 3.1(3), we have
$\overline{f}\leq\phi^{\text{cd}}_{\xi}(x)\leq\overline{f}.$
Therefore, we have
@@
Suppose Assumptions 3.2 holds. We denote by
@@
where $r^{\prime}_{y_{1}}(x,y_{1})=\mathcal{O}(y_{1}^{3})$ with respect to $y_{1}$. It is clearly that under this new coordinate, the stationary point must satisfies $y_{2}=\cdots=y_{m}=0$. We next eliminate the linear term in the first equation of (C.10) to facilitate the analysis of the solution structure of (C.10). Consider the following map:
@@
$F(0,0)=b(0)=0,\quad\frac{\partial F}{\partial y_{1}}(0,0)=c(0)\neq 0,$
by the implicit function theorem, we obtain that there exists a continuously differentiable function $Y_{1}(x)$ satsifies $Y_{1}(0)=0$ and $F(x,Y_{1}(x))=0$. Therefore, the first equation of (C.10) can be written as
@@
$A(x)$ $=a(x)+b(x)Y_{1}(x)+c(x)(Y_{1}(x))^{2}+r^{\prime}_{y_{1}}(x,Y_{1}(x))$
$C(x)$ $=2c(x)+r^{\prime\prime\prime}_{y_{1}y_{1}y_{1}}(x,Y_{1}(x)).$
It is easy to see that $A(x)$ and $C(x)$ satisfy $A(0)=0$, $C(0)\neq 0$, and
@@
$\alpha<C(x)+\frac{R(x,y_{1})}{y_{1}^{2}}<\beta\quad\text{or}\quad-\beta<C(x)+\frac{R(x,y_{1})}{y_{1}^{2}}<-\alpha.$
The specific case depends on the sign of $C(0)$. Without loss of generality, we assume that $C(0)>0$, i.e., we are in the first case. From (C.15) we can see
@@
$G(x,\sqrt{\frac{-A(x)}{\alpha}})>A(x)-\frac{A(x)}{\alpha}\cdot\alpha=0$
$G(x,\sqrt{\frac{-A(x)}{\beta}})<A(x)-\frac{A(x)}{\beta}\cdot\beta=0$
$G(x,-\sqrt{\frac{-A(x)}{\alpha}})>A(x)-\frac{A(x)}{\alpha}\cdot\alpha=0$
$G(x,-\sqrt{\frac{-A(x)}{\beta}})<A(x)-\frac{A(x)}{\beta}\cdot\beta=0.$
@@
We proceed to prove that these coordinate transformations preserve the stationary points, and then investigate bounds on $|\lambda(x,y)|$ in the original coordinates. We denote the original coordinate system by $(x,y)$, the new coordinate system by $(x^{\prime},y^{\prime})$. Note that we in fact applied only a coordinate transformation on the $x$-component, independent of $y$. Therefore, the coordinate transformation can be written as follows:
@@
which is invertible. By the inverse function theorem, this mapping defines a local coordinate transformation, and is therefore bi-Lipschitz. Without loss of generality, we may assume this holds in $W$. In the new coordinate system, the set $\mathsf{proj}_{x}(\Gamma\cap W)$ corresponds exactly to the set of points where $x_{1}^{\prime}=0$. A point $(x_{1},...,x_{n})$ in the original coordinates maps to $(x_{1}^{\prime},...,x_{n}^{\prime})$ in the new coordinates, where $x_{1}^{\prime}=A(x)$. Since the transformation is bi-Lipschitz, the distance from $(x_{1},...,x_{n})$ to $\mathsf{proj}_{x}(\Gamma\cap W)$ in the original coordinates is upper bounded by $L|A(x)|$ for some constant $L>0$. Combining with (C.20), we obtain that there exists two constants $C_{1}$ and $C_{2}$ such that the following holds for any stationary point $(x,y)$ in $W$
@@
Note that the rate of $\alpha(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ with respect to $\delta$ is exactly the rate, in terms of $\delta$, of the minimal gradient norm of $g$ with respect to $y$ over the set $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$. Our approach is as follows: for each stationary point in $\mathcal{X}\times\mathcal{Y}$, we construct an open neighborhood $W$ that is independent of $\delta$. Since the set of stationary points in $\mathcal{X}\times\mathcal{Y}$ is closed, we can find finitely many such open neighborhoods $\{W_{i}\}_{i=1}^{I}$ that cover the entire set of stationary points. Then, on $(\mathcal{X}\times\mathcal{Y})\setminus(\cup_{i=1}^{I}W_{i})$, the gradient norm of $g$ with respect to $y$ naturally admits a positive lower bound independent of $\delta$. Therefore, it suffices to analyze the lower bound of the gradient norm within each set $W_{i}\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$.
We construct open neighborhoods separately according to the following cases:
  1. 1.
    The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\notin\widetilde{\mathcal{X}}$. This means that $\nabla_{yy}^{2}g(\bar{x},\bar{y})$ is non-degenerate. By the implicit function theorem, $\bar{y}$ can be locally extended to a continuous function $\bar{y}(x)$ for $x\in B_{\bar{x}}(\Delta_{x})$, where $B_{\bar{x}}(\Delta_{x})$ denotes the ball centered at $\bar{x}$ with radius $\Delta_{x}$. Here $\Delta_{x}$ is chosen sufficiently small, less than half of the distance from $\bar{x}$ to the closed set $\widetilde{\mathcal{X}}$. Therefore, $\nabla_{yy}^{2}g(x,\bar{y}(x))$ remains non-degenerate for all $x\in B_{\bar{x}}(\Delta_{x})$, and the absolute values of all eigenvalues of $\nabla_{yy}^{2}g(x,\bar{y}(x))$ admit a uniform positive lower bound, denoted by $\lambda$. We then define the open set corresponding to $(\bar{x},\bar{y})$ as
@@
    Next, we consider the intersection of $W$ and $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$, and estimate the lower bound of the gradient norm of $g$ with respect to $y$ on this intersection. To ensure that the intersection is non-empty, we may assume $\delta<\Delta_{x}$ and $\mu(\delta)<\lambda$. Fix $x\in B_{\bar{x}}(\Delta_{x})$, and consider the slice $\{x\}\times B_{\bar{y}(x)}(\lambda/(2\overline{\overline{L}}_{g}))$. By Lipschitz continuity of Hessian matrix (from Assumption 3.2(4)), we know that for $y\in B_{\bar{y}(x)}(\lambda/2\overline{\overline{L}}_{g})$, it holds that
@@
  2. 2.
    The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\in\widetilde{\mathcal{X}}$, but it is a non-degenerate stationary point, i.e., $\nabla_{yy}g(\bar{x},\bar{y})$ is non-degenerate. This case is almost the same as Case 1. Since $(\bar{x},\bar{y})$ is a non-degenerate stationary point, we can also construct a locally defined continuous curve of stationary points $\bar{y}(x)$. Along this curve, the absolute values of the eigenvalues of the Hessian admit a uniform positive lower bound $\lambda$, and thus we can construct $W$ in the same way as in Case 1. The only difference is that now the intersection of $W$ and $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ is further intersected with $(\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta})\times\mathcal{Y}$ compared to Case 1. However, taking this additional intersection can only increase the minimal value of the gradient norm, and hence, we still obtain
@@
  3. 3.
    The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\in\widetilde{\mathcal{X}}$, and it is a degenerate stationary point, i.e., $\nabla_{yy}g(\bar{x},\bar{y})$ is degenerate. We first perform the coordinate transformation (C.18) in the proof of Lemma C.1. In this coordinate, the equation system of stationary points has the form (C.14). Recall the functions $A$, $C$, $R$ in (C.14), we suppose that $C(0)>0$. We further perform the following two coordinate transformations
@@
    We denote the original coordinate system by $(x,y)$, and the new system after performing the coordinate transformation (C.18) and $\Phi_{1}$, $\Phi_{2}$ by $(x^{\prime},y^{\prime})$. The coordinate mappings is denoted by $\Psi:(x,y)\mapsto(x^{\prime},y^{\prime})$. Under this new coordinate, the equation system of stationary points has the form:
@@
    Without loss of generalty, we suppose $(\bar{x},\bar{y})=(0,0)$. It is clearly that $(\bar{x}^{\prime},\bar{y}^{\prime})=\Psi(x,y)=(0,0)$. We then choose the open neighborhood $W$ of $(\bar{x},\bar{y})=(0,0)$ such that $\Psi(W)$ is a box neighborhood of $(\bar{x}^{\prime},\bar{y}^{\prime})=(0,0)$ of the form $\{(x^{\prime},y^{\prime}):|x^{\prime}_{i}|<r,\ |y^{\prime}_{j}|<r\text{ for all }i,j\}$ for some constant $r>0$, i.e.,

### v3

@@ Appendix C Detailed proof
From Nesterov and Polyak (2006, Theorem 1), for the first $K_{1}$ iterations, the cubic-regularized Newton method guarantees:
$\min_{k=1,\dots,K_{1}}\nu_{\overline{\overline{L}}g}(y_{k})\leq\frac{8}{3}\left(\frac{3(g(x,y_{0})-\underline{g})}{2K_{1}\overline{\overline{L}}g}\right)^{1/3},$ (C.1)
where the stationarity measure is defined as
$\nu_{M}(y):=\max\left\{\sqrt{\frac{1}{M}|\nabla_{y}g(x,y)|},-\frac{2}{3M}\lambda{\min}(\nabla^{2}_{yy}g(x,y))\right\}.$
Let $k^{*}$ be the index minimizing this measure within the first $K_{1}$ steps. By the specific choice of $K_{1}$ in (4.4), substituting $K_{1}$ into the RHS of (C.1) ensures that $y_{k^{*}}$ satisfies two key conditions:
1. Gradient Condition: The gradient norm satisfies:
$|\nabla_{y}g(x,y_{k^{*}})|<\min\left\{\alpha\left(\delta,\frac{\mu(\delta)}{2\overline{\overline{L}}_{g}}\right),\frac{(\mu(\delta))^{2}}{4\overline{\overline{L}}_{g}}\right\}.$ (C.2)
2. Curvature Condition: The smallest eigenvalue of the Hessian satisfies:
$\lambda_{\min}(\nabla^{2}_{yy}g(x,y_{k^{*}}))>-\frac{\mu(\delta)}{2}.$ (C.3)
We now analyze the geometric implications of these bounds.
Phase I: Localization to a Strongly Convex Neighborhood. Since $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$, $g(x,\cdot)$ is Morse. Let $\{p_{j}\}$ denote its discrete stationary points in $\mathcal{Y}^{\prime}$. From Lemma 4.3, the first term in the minimum of (C.2) implies that $y_{k^{*}}$ must lie within the $r$-neighborhood of some stationary point $p_{j}$, where $r:=\mu(\delta)/(2\overline{\overline{L}}_{g})$. Specifically, $y_{k^{*}}\in\bigcup_{j}B(p_{j},r)$. Within any such ball $B(p_{j},r)$, the Lipschitz continuity of the Hessian (Assumption 3.2(4)) and Lemma 4.2 imply that all eigenvalues of $\nabla^{2}_{yy}g(x,y)$ are lower bounded by $\mu(\delta)-\overline{\overline{L}}_{g}\cdot r=\mu(\delta)/2$ in absolute value. However, for saddle points or local maximizers, the smallest eigenvalue would be less then or equal to $\mu(\delta)/2$. Condition (C.3) ($\lambda_{\min}>-\mu(\delta)/2$) contradicts this, thereby excluding balls centered at non-minimizers. Furthermore, the monotonicity of the algorithm (Assumption 3.3) excludes any local minimizers outside $\mathcal{Y}$ that have function values higher than $g(x,y_{0})$. Thus, $y_{k^{*}}$ must reside in $B(p_{j^{*}},r)$ for some local minimizer $p_{j^{*}}\in\mathcal{Y}$, denoted hereafter as $y^{\text{cd}}(x)$. In this neighborhood $B(y^{\text{cd}}(x),r)$, $g(x,\cdot)$ is strongly convex with modulus $\mu(\delta)/2$. The second term in (C.2), $\|\nabla_{y}g(x,y_{k^{*}})\|<(\mu(\delta))^{2}/(4\overline{\overline{L}}_{g})$, ensures that $y_{k^{*}}$ lies the ball $B(p_{j^{*}},r)$. Specifically, it belongs to a local sublevel set $\mathcal{F}=\{y\in B(y^{\text{cd}}(x),r):g(x,y)-g(x,y^{\text{cd}}(x))\leq(\mu(\delta))^{3}/(16\overline{\overline{L}}_{g}^{2})\}$. By Nesterov and Polyak (2006, Lemma 2), the algorithm will never leave this region $\mathcal{F}$ in subsequent iterations.
Phase II: Final Convergence. The sequence now evolves inside a region where $g(x,\cdot)$ is strongly convex. Starting from $y_{k^{*}}$, we perform an additional $K_{2}$ iterations. Applying Nesterov and Polyak (2006, Theorem 1) again for these $K_{2}$ steps, the choice of $K_{2}$ in (4.4) guarantees finding a point $\hat{y}(x)$ such that:
$|\nabla_{y}g(x,\hat{y}(x))|\leq\frac{\mu(\delta)\rho}{2}.$
Finally, using the $\frac{\mu(\delta)}{2}$-strong convexity of $g(x,\cdot)$ in the ball $B(y^{\text{cd}}(x),r)$, we have:
$|\hat{y}(x)-y^{\text{cd}}(x)|\leq\frac{1}{\mu(\delta)/2}|\nabla_{y}g(x,\hat{y}(x))|\leq\frac{2}{\mu(\delta)}\cdot\frac{\mu(\delta)\rho}{2}=\rho.$
This completes the proof.
@@ 2. Curvature Condition: The smallest eigenvalue of the Hessian satisfies:
Note that $K$ satisfies (4.4), according to Lemma 4.4, we have $\|\hat{y}(z)-y^{\text{cd}}(z)\|\leq\rho$ for any $z\in\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta}$. By Lipschitz continuity of $f$ (from Assumption 3.2(7)), it follows $|f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))|\leq L_{f}\rho$. From (3.8), we know that $|f(z,y^{\text{cd}}(z))|\leq\overline{f}$ for any $z\in\mathcal{X}$. By an argument similar to that of (3.8), we have $|f(z,\hat{y}(z))|\leq\overline{f}$ for any $z\in\mathcal{X}$ as well. Thus, the following holds:
@@
Note that we assume that the number of lower-level iterations satisfies $K\geq K_{0}$ in Assumption 3.3. $y^{\text{cd}}(x)$ is always in the level set $\{y:g(x,y)\leq g(x,y_{0})\}$. By Proposition 3.1(2), we also have $\overline{f}\leq\phi^{\text{cd}}_{\xi}(x)\leq\overline{f}.$ It follows
@@
Suppose $g(x,y)$ is three times continuously differentiable, and Assumptions 3.2 holds. We denote by
@@
where $r^{\prime}_{y_{1}}(x,y_{1})=\mathcal{O}(y_{1}^{3})$ with respect to $y_{1}$. It is clear that under this new coordinate, the stationary point must satisfy $y_{2}=\cdots=y_{m}=0$. We next eliminate the linear term in the first equation of (C.10) to analyze the solution structure of (C.10). Consider the following map:
@@
$F(0,0)=b(0)=0,\quad\frac{\partial F}{\partial y_{1}}(0,0)=2c(0)\neq 0,$
by the implicit function theorem, we obtain that there exists a continuously differentiable function $Y_{1}(x)$ satisfies $Y_{1}(0)=0$ and $F(x,Y_{1}(x))=0$. Therefore, the first equation of (C.10) can be written as
@@
$A(x)$ $=\frac{\partial g}{\partial y_{1}}(x,Y_{1}(x))=a(x)+b(x)Y_{1}(x)+c(x)(Y_{1}(x))^{2}+r^{\prime}_{y_{1}}(x,Y_{1}(x))$
$C(x)$ $=\frac{1}{2}\frac{\partial^{3}g}{\partial y_{1}^{3}}(x,Y_{1}(x))=c(x)+\frac{1}{2}r^{\prime\prime\prime}_{y_{1}y_{1}y_{1}}(x,Y_{1}(x)).$
It is easy to see that $A(x)$ and $C(x)$ satisfy $A(0)=a(0)=0$, $C(0)=c(0)\neq 0$, and
@@
$\alpha<C(x)+\frac{R(x,y_{1})}{y_{1}^{2}}<\beta\quad\text{or}\quad-\beta<C(x)+\frac{R(x,y_{1})}{y_{1}^{2}}<-\alpha.$ (C.16)
The specific case depends on the sign of $C(0)$. Without loss of generality, we assume that $C(0)>0$, i.e., we are in the first case of (C.16). From (C.15) we can see
@@
$G(x,\sqrt{\frac{-A(x)}{\alpha}})>A(x)-\frac{A(x)}{\alpha}\cdot\alpha=0\quad G(x,\sqrt{\frac{-A(x)}{\beta}})<A(x)-\frac{A(x)}{\beta}\cdot\beta=0$
$G(x,-\sqrt{\frac{-A(x)}{\alpha}})>A(x)-\frac{A(x)}{\alpha}\cdot\alpha=0\quad G(x,-\sqrt{\frac{-A(x)}{\beta}})<A(x)-\frac{A(x)}{\beta}\cdot\beta=0.$
@@
We proceed to prove that these coordinate transformations preserve the stationary points, and then investigate bounds on $|\lambda(x,y)|$ in the original coordinates. We denote the original coordinate system by $(x,y)$, the new coordinate system by $(x^{\prime},y^{\prime})$. Note that, for the $x$-component, we only applied a coordinate transformation independent of $y$. Therefore, the coordinate transformation can be written as follows:
@@
which is invertible. By the inverse function theorem, this mapping defines a local coordinate transformation, and is therefore bi-Lipschitz. By shrinking $W$ if necessary, we may assume that the mapping is bi-Lipschitz on the entire set $W$. In the new coordinate system, the set $\mathsf{proj}_{x}(\Gamma\cap W)$ corresponds exactly to the hyperplane $x_{1}^{\prime}=0$. Consequently, the distance from a point $x^{\prime}$ to this set is simply $|x_{1}^{\prime}|=|A(x)|$. Invoking the bi-Lipschitz property, the distance in the original coordinates is bounded by $\mathrm{dist}(x,\text{Proj}_{x}(\Gamma\cap W))\leq L|A(x)|$ for some constant $L>0$. Combining with (C.21), we obtain that there exists two constants $C_{1}$ and $C_{2}$ such that the following holds for any stationary point $(x,y)$ in $W$
@@
Note that the rate of $\alpha(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ with respect to $\delta$ is exactly the rate, in terms of $\delta$, of the minimal gradient norm of $g$ with respect to $y$ over the set $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$. Our approach is as follows: for each stationary point in $\mathcal{X}\times\mathcal{Y}$, we construct an open neighborhood $W$ that is independent of $\delta$. Since the set of stationary points in $\mathcal{X}\times\mathcal{Y}$ is closed, we can find finitely many such open neighborhoods $\{W_{i}\}_{i=1}^{I}$ that cover the entire set of stationary points. Then, on $(\mathcal{X}\times\mathcal{Y})\setminus(\cup_{i=1}^{I}W_{i})$, the gradient norm of $g$ with respect to $y$ naturally admits a positive lower bound independent of $\delta$. Therefore, it suffices to analyze the lower bound of the gradient norm within each set $W_{i}\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$. We construct open neighborhoods separately according to the following cases:
The first case: The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\notin\widetilde{\mathcal{X}}$. This means that $\nabla_{yy}^{2}g(\bar{x},\bar{y})$ is non-degenerate. By the implicit function theorem, $\bar{y}$ can be locally extended to a continuous function $\bar{y}(x)$ for $x\in B_{\bar{x}}(\Delta_{x})$, where $B_{\bar{x}}(\Delta_{x})$ denotes the ball centered at $\bar{x}$ with radius $\Delta_{x}$. Here $\Delta_{x}$ is chosen sufficiently small such that $\nabla_{yy}^{2}g(x,\bar{y}(x))$ remains non-degenerate for all $x\in B_{\bar{x}}(\Delta_{x})$. Then the absolute values of all eigenvalues of $\nabla_{yy}^{2}g(x,\bar{y}(x))$ admit a uniform positive lower bound, denoted by $\lambda$. We then define the open set corresponding to $(\bar{x},\bar{y})$ as
@@
Next, we consider the intersection of $W$ and $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$, and estimate the lower bound of the gradient norm of $g$ with respect to $y$ on this intersection. To ensure that the intersection is non-empty, we may assume $\delta<\Delta_{x}$ and $\mu(\delta)/(2\overline{\overline{L}}_{g}))<\lambda/(2\overline{\overline{L}}_{g}))$. Fix $x\in B_{\bar{x}}(\Delta_{x})$, and consider the slice $\{x\}\times B_{\bar{y}(x)}(\lambda/(2\overline{\overline{L}}_{g}))$. By Lipschitz continuity of Hessian matrix (from Assumption 3.2(4)), we know that for $y\in B_{\bar{y}(x)}(\lambda/2\overline{\overline{L}}_{g})$, it holds that
@@
The second case: The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\in\widetilde{\mathcal{X}}$, but it is a non-degenerate stationary point, i.e., $\nabla_{yy}g(\bar{x},\bar{y})$ is non-degenerate. This case is almost the same as Case 1. Since $(\bar{x},\bar{y})$ is a non-degenerate stationary point, we can also construct a locally defined continuous curve of stationary points $\bar{y}(x)$. Along this curve, the absolute values of the eigenvalues of the Hessian admit a uniform positive lower bound $\lambda$, and thus we can construct $W$ in the same way as in Case 1. The only difference is that now the intersection of $W$ and $K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ is further intersected with $(\mathcal{X}\setminus\widetilde{\mathcal{X}}_{\delta})\times\mathcal{Y}$ compared to Case 1. However, taking this additional intersection can only increase the minimal value of the gradient norm, and hence, we still obtain
@@
The third case: The pair $(\bar{x},\bar{y})$ is a stationary point with $\bar{x}\in\widetilde{\mathcal{X}}$, and it is a degenerate stationary point, i.e., $\nabla_{yy}g(\bar{x},\bar{y})$ is degenerate. We first perform the coordinate transformation (C.19) in the proof of Lemma C.2. In this coordinate, the equation system of stationary points has the form (C.14). Recall the functions $A$, $C$, $R$ established in (C.14), we suppose that $C(0)>0$. We further perform the following two coordinate transformations
@@
We denote the original coordinate system by $(x,y)$, and the new system after performing the coordinate transformation (C.19) and $\Phi_{1}$, $\Phi_{2}$ by $(x^{\prime},y^{\prime})$. The coordinate mapping is denoted by $\Psi:(x,y)\mapsto(x^{\prime},y^{\prime})$. Under this new coordinate, the equation system of stationary points has the form:
@@
Without loss of generality, we suppose $(\bar{x},\bar{y})=(0,0)$. It is clear that $(\bar{x}^{\prime},\bar{y}^{\prime})=\Psi(x,y)=(0,0)$. We then choose the open neighborhood $W$ of $(\bar{x},\bar{y})=(0,0)$ such that $\Psi(W)$ is a box neighborhood of $(\bar{x}^{\prime},\bar{y}^{\prime})=(0,0)$ of the form $\{(x^{\prime},y^{\prime}):|x^{\prime}_{i}|<r,\ |y^{\prime}_{j}|<r\text{ for all }i,j\}$ for some constant $r>0$, i.e.,